How to use substitution?
X=-4Y
and
3X+2Y=20
Im lost..please help me!

Answer by rfer(16322)

3(-4y)+2y=20
-12y+2y=20
-10y=20
y=-2
---------------
3x+2*-2=20
3x-4=20
3x=24
x=8
--------------
proof
x=-4y
8=-4*-2